OpenStudy (anonymous):

The slope of a line is 1/8. The point (8,9) is on the line. Determine the equation of this line

OpenStudy (ivettef365):

you need to use the equation y = mx + b

OpenStudy (ivettef365):

do you know how to use the equation

OpenStudy (anonymous):

No, My teacher never taught me it, she just gave me the assignment

OpenStudy (ivettef365):

oh great, well let me explain, on the equation y=mx+b y= a point on the line x= a point on the line m= slope b = y-intercept

OpenStudy (ivettef365):

now you have slope of 1/8, points (8,9) equivalent to (x,y) all you need to find is the y-int, value of b

OpenStudy (ivettef365):

so the way you set up the equation as follows: y = mx + b 9 = 1/8(8) + b can you solve for b

OpenStudy (anonymous):

is b 8?

OpenStudy (ivettef365):

correct

OpenStudy (ivettef365):

so the way you right the equation as follows y = 1/8x + 8 (you don't put values on x,y, because that represents all the point that passes through that line

OpenStudy (anonymous):

ohh okay thank you

OpenStudy (ivettef365):

yw :)